OpenStudy (anonymous):

Prove that dx/dy =1/dy/dx .

OpenStudy (solomonzelman):

\(\large\color{blue}{ \frac{a}{b} \div \frac{c}{d}= \frac{a}{b} \times \frac{d}{c} }\) Use this rule to simplify the right side.

OpenStudy (anonymous):

I'm not too sure about that. This is one of the questions i got for my calc exam. Also proofing questions are never one line long.

OpenStudy (anonymous):

The question is pretty much asking to prove that a function is equal to the reciprocal of the inverse function/

OpenStudy (masumanwar):

dx/dy is the definition for derivative of x with respect to y if we now change the derivative with respect to x for y then we should divide with dx in denominator and numerator to get 1/dy/dx

OpenStudy (anonymous):

http://www.physicsforums.com/showthread.php?t=63886